OpenStudy (anonymous):

I need help with math, heres my problem 1/2 (6x - 4) = 4x -9

OpenStudy (anonymous):

can anyone please help me?????

OpenStudy (anonymous):

multiply 2 on both sides 6x - 4 = 8x - 18 -2x = -14 x = 7 cheers

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

also i need help with 15-x=2 (x+3)

OpenStudy (anonymous):

15 - x = 2x + 6 x = 9

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

I DID IT WRONG!! i distributed the 2 and got 15 - x = 2x + 6 add x on both sides 15 = 3x + 6 subtract 6 on both sides 3x = 9 divide 3 on both sides x = 3 it's actually 3 sorry

OpenStudy (anonymous):

ok thanks